A CORRECTION TO “BASICS OF O-MINIMALITY AND HARDY FIELDS”

CHRIS MILLER

Version: December 1, 2018

This is not a preprint; please do not refer to it as such.

It has been noticed (by P. Speissegger) that the leftmost = in the last display of [1, p. 57] should be <.

References

[1] Chris Miller, Basics of o-minimality and Hardy fields, Lecture notes on o-minimal structures and real analytic geometry, Fields Inst. Commun., vol. 62, Springer, New York, 2012, pp. 43– 69. MR 2976990

Department of Mathematics, The Ohio State University, 231 West 18th Avenue, Columbus, Ohio 43210, USA

E-mail address: miller@math.osu.edu